65 miles per hour is equal to how many feet per second?
Answer by checkley75(3666)

65 MILES=65*5,280=343,200 FEET
1 HOUR=3,600 SECONDS
NOW WE DIVIDE THESE TWO TERMS
343,200/3600=95.3333 FEET PER SECOND.